OpenStudy (anonymous):

Maria flips a dime four times. What is the probability that she will get at least one head?

OpenStudy (anonymous):

I know for sure it is 1/2 Would I multiply 1/2 4 times?

sammixboo (sammixboo):

How many sides are on a coin?

OpenStudy (anonymous):

2.

sammixboo (sammixboo):

That means she has a 50% chance every time she flips a head

sammixboo (sammixboo):

If she flips a coin 4 times she still has a 50-50 chance of getting a head

OpenStudy (anonymous):

So it would be 1/2*1/2*1/2*1/2 =1/16? Is that right

OpenStudy (ikram002p):

at least one head means one head and one write + two head no write 1/2*1/2 +1/2*1/2=2/4=1/2

ganeshie8 (ganeshie8):

you could also find the probability of "no heads" and subtract it from 1

OpenStudy (ikram002p):

yeah hehehe

OpenStudy (anonymous):

\[1-\left(\frac{1}{2}\right)^4=\frac{15}{16} \]